# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

### 16 Mar 2026 v4

@@ Abstract
the polarization acquires a universal singular form
$\Pi(\omega)\sim\frac{\rho_{0}}{|\omega|}.$ (38)
In the time domain this corresponds to a long-memory kernel
$K(t)\sim\frac{1}{t}.$ (39)
Importantly, this singular behavior does not originate from Landau damping into a fast electronic bath. Instead it arises from the spectral accumulation of slow relaxation modes themselves. A complete derivation of Eq. (36) from the stochastic dynamics of relaxational pair modes is presented in Appendix D.
III.3 C. Infrared amplification of pairing correlations
In conventional superconductivity theories the transition temperature is often obtained from a microscopic gap equation derived from fermionic quasiparticle propagators. Such formulations rely on the existence of reasonably well-defined single-particle excitations, so that the Cooper kernel can be written in terms of products of fermionic Green functions.
In strongly correlated systems, however, the quasiparticle picture may cease to be valid. The single-particle spectral function can become broad and incoherent, and the low-energy dynamics may instead be governed by a continuum of collective relaxation processes. Within the present framework the infrared dynamics is therefore organized in terms of the relaxation-rate spectrum $\rho(\lambda)$ rather than quasiparticle propagation.
Because the microscopic quasiparticle propagator is not introduced explicitly, the superconducting transition is identified through the instability of the Cooper-channel susceptibility. The corresponding Thouless criterion reads
$\chi_{\Delta}^{-1}(0,T_{c})=0,\quad\text{or equivalently}\quad 1-g_{\rm pair}\Pi(T_{c})=0.$ (40)
Here $g_{\rm pair}$ represents the intrinsic electronic interaction that seeds Cooper correlations, while the relaxation spectrum controls their dynamical amplification through the infrared structure of $\Pi(\omega)$.
The Cooper susceptibility therefore takes the form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{pair}}-\Pi(\omega).$ (41)
Because $\Pi(\omega)$ diverges as $|\omega|\to 0$, pairing correlations are strongly enhanced as the energy scale is lowered.
To understand the infrared flow more generally, consider a TDOS scaling
$\rho(\lambda)\sim\lambda^{\eta}.$ (42)
Substituting this form into the polarization integral yields
$\Pi(\omega)\sim\int d\lambda\frac{\lambda^{\eta}}{\lambda^{2}+\omega^{2}}\sim|\omega|^{-(1-\eta)}.$ (43)
The Cooper-channel kernel therefore acquires the infrared scaling
$\Pi(\omega)\sim|\omega|^{\eta-1}.$ (44)
Under RG rescaling $\omega\rightarrow\omega/b$ the effective pairing interaction transforms as
$g^{\prime}=b^{1-\eta}g.$ (45)
Introducing $l=\ln b$ gives the tree–level RG flow
$\frac{dg}{dl}=(1-\eta)g.$ (46)
For $\eta<1$ the Cooper interaction becomes relevant in the infrared. Including the usual ladder correction yields
$\frac{dg}{dl}=(1-\eta)g+Cg^{2}.$ (47)
The linear term dominates the early RG flow, leading to algebraic growth
$g(l)=g_{0}e^{(1-\eta)l}.$ (48)
Physical interpretation. The qualitative difference between the Eliashberg regime and the present memory–dominated regime originates from the structure of the collective spectrum, as illustrated schematically in Fig. 2.
In the Eliashberg case, each logarithmic shell contributes independently, leading to marginal enhancement of pairing correlations. In contrast, a flat TDOS corresponds to a dense continuum of near–marginal relaxation modes. Fluctuations across widely separated time scales remain dynamically correlated, allowing pairing correlations to accumulate coherently.
Superconductivity therefore emerges as a dynamical instability of the collective relaxation spectrum itself, rather than from the exchange of a specific bosonic mediator.
Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction. Shown schematically is the flow of the dimensionless pairing coupling $g$ under infrared coarse graining. Ohmic Landau damping of collective modes produces marginal logarithmic growth, $dg/dl\sim g^{2}$, characteristic of conventional Eliashberg theory (blue). In contrast, memory–dominated dynamics generated by a flat relaxation–rate density of states yields a relevant infrared instability, $dg/dl=(1-\eta)g$, leading to algebraic amplification of the pairing interaction (orange). The slow–mode reservoir therefore dynamically enhances intrinsic electronic pairing tendencies rather than merely renormalizing a fixed microscopic coupling.
IV IV. Pairing from Memory–Dominated Critical Dynamics
The dynamical framework developed in the preceding sections establishes that strongly correlated matter generically self–organizes into a regime dominated by an extensive reservoir of slow collective modes. Rather than producing conventional overdamped fluctuations described by Ohmic Eliashberg dynamics, the accumulation of near–marginal relaxation rates generates long–time memory kernels and nonanalytic infrared response.
Superconductivity in this setting emerges as a dynamical instability of the memory–dominated collective spectrum. The Cooper channel is reorganized by the slow–mode reservoir into a relevant infrared interaction, fundamentally distinct from marginal pairing mediated by fast bosonic modes. Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into an algebraic superconducting instability.
In this section we demonstrate explicitly how microscopic electronic correlations seed pairing and how the slow–mode TDOS reorganizes the Cooper channel into a memory–enhanced critical regime, ultimately yielding algebraic scaling of the transition temperature and robust phase coherence.
IV.1 A. Intrinsic electronic pairing and dynamical formulation
In strongly correlated materials proximate to Mott insulating phases, short–range Coulomb interactions naturally generate local singlet correlations through exchange processes [1, 30, 31]. Starting from the Hubbard Hamiltonian,
$H=-t\sum_{\langle ij\rangle\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}+U\sum_{i}n_{i\uparrow}n_{i\downarrow},$ (49)
projection to the strong–coupling sector yields the superexchange interaction
$H_{J}=J\sum_{\langle ij\rangle}\Big(\mathbf{S}_{i}\cdot\mathbf{S}_{j}-\frac{1}{4}n_{i}n_{j}\Big),\qquad J=\frac{4t^{2}}{U},$ (50)
which favors local singlet formation. This interaction provides a natural microscopic origin of electronic pairing in systems near Mott localization.
More generally, interacting electrons are described by a generic four–fermion interaction
$H_{\rm int}=\sum_{k,k^{\prime},q}V(q)\,c^{\dagger}_{k+q\sigma}c^{\dagger}_{k^{\prime}-q\sigma^{\prime}}c_{k^{\prime}\sigma^{\prime}}c_{k\sigma}.$ (51)
This interaction contains multiple collective channels. In conventional diagrammatic approaches these channels are analyzed through resummations such as particle–hole bubble diagrams or particle–particle ladder diagrams. Rather than focusing on a specific mediator, the present framework emphasizes the dynamical organization of the collective spectrum itself.
While the four–fermion interaction generically allows multiple ordering channels, their ordering tendencies are not equivalent. Particle–hole instabilities such as charge or spin density waves typically require specific momentum structure, most notably Fermi surface nesting or commensurate wave vectors determined by the lattice. Such kinematic constraints restrict the available phase space for density–wave ordering.
Superconductivity, by contrast, corresponds to condensation in the particle–particle channel at zero total momentum,
$\Delta=\langle c_{k\uparrow}c_{-k\downarrow}\rangle,$ (52)
and therefore does not rely on special nesting conditions of the Fermi surface. Consequently the Cooper channel can efficiently accumulate infrared spectral weight generated by collective fluctuations.
Within the dynamical framework developed in Sec. III, the pairing susceptibility assumes the universal tt–matrix form in Eq. (41), where $g_{\rm pair}$ characterizes the intrinsic electronic pairing scale $E_{\rm pair}$ generated by short–range correlations. Diagrammatically this corresponds to the Cooper-channel polarization obtained after integrating out microscopic electronic degrees of freedom.
In weakly interacting systems the pairing instability is exponentially suppressed as in conventional BCS theory. Near correlated criticality, however, the situation changes qualitatively. Because the relaxation spectrum becomes dense near $\lambda\to 0$, an extensive reservoir of slow collective modes emerges. This slow–mode reservoir reorganizes the infrared dynamics of the electronic system and strongly enhances the Cooper channel.
Within the present framework critical dynamics does not create attraction. Instead it amplifies a preexisting electronic pairing tendency through infrared spectral reorganization. The intrinsic pairing scale $E_{\rm pair}$ therefore acts as the microscopic seed, while the slow collective spectrum dynamically amplifies the Cooper instability.
IV.2 B. Memory–dominated Cooper kernel from flat TDOS
Because the collective pair fluctuations inherit the relaxational spectral structure generated by the slow-mode reservoir, the Cooper-channel polarization kernel reduces to a continuum integral over relaxation rates,
$\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\rho_{\rm eff}(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (53)
When the slow–mode reservoir develops,
$\rho_{\rm eff}(\lambda\to 0)=\rho_{0}\neq 0,$ (54)
the infrared behavior becomes
$\Pi(\omega)\xrightarrow{|\omega|\ll\Lambda}\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|}.$ (55)
This nonanalytic enhancement reflects the accumulation of collective memory across widely separated time scales. Physically, the flat TDOS implies that an extensive set of slow collective modes contributes coherently to the Cooper channel. Unlike conventional Eliashberg kernels, which remain marginal in the infrared, the present Cooper kernel becomes strongly enhanced by the infrared spectral weight of the slow–mode reservoir. The pairing interaction is therefore dynamically amplified by the collective memory of the system.
IV.3 C. Algebraic instability and transition scale
At finite temperature the imaginary-time direction becomes compact, $\tau\in[0,\beta]$ with $\beta=1/T$. As a consequence the system cannot probe arbitrarily long time scales. In the imaginary-time formalism the dynamical fields are expanded in bosonic Matsubara frequencies
$\Omega_{n}=2\pi nT.$ (56)
The smallest nonzero dynamical frequency therefore sets a natural infrared scale of order
$|\omega|\sim T.$ (57)
Substituting this scale into the infrared expression (55) gives
$\Pi(T)\sim\frac{\rho_{0}}{T}.$ (58)
Thus the infrared divergence present at zero temperature is universally regulated by the thermal scale.
The superconducting transition follows from the Thouless criterion
$\chi_{\Delta}^{-1}(\omega\sim T_{c})=0,$ (59)
which yields the algebraic transition scale
$T_{c}\sim\rho_{0}\,E_{\rm pair},$ (60)
up to numerical factors of order unity.
The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling governed by the infrared spectral weight of the slow–mode reservoir. High transition temperatures arise generically from dynamical amplification of moderate intrinsic pairing scales.
It is important to emphasize that superconductivity requires not only pair formation but also global phase coherence. The transition temperature is therefore ultimately bounded by the smaller of the pairing amplitude and the superfluid phase stiffness,
@@
with $\rho_{s}\sim n_{s}/m^{*}$. The detailed scaling of the phase stiffness within the slow–mode framework is discussed in Sec. V.
The key result of the present section is that the pairing scale itself is controlled by a single infrared parameter, the zero–frequency TDOS $\rho_{0}$. The slow–mode reservoir therefore acts as a universal dynamical amplifier: it converts intrinsic electronic correlation energy into an algebraically enhanced instability of the Cooper channel, independent of any specific bosonic mediator.
In this sense, the TDOS–controlled fixed point defines a dynamical stability class distinct from both weak–coupling BCS superconductors and conventional Eliashberg scenarios. The emergence of superconductivity is governed not by the exchange of a particular glue, but by infrared spectral organization of collective relaxation modes.
Figure 3: Experimental consequences of memory–dominated critical dynamics. (a) Superconducting dome emerging from truncation of the slow–mode reservoir away from criticality. The transition temperature follows $T_{c}(p)\sim E_{\rm pair}\rho_{0}(p)$ and is maximal near the correlated critical point $p_{c}$, where infrared spectral weight of collective relaxation modes is largest. Suppression of $T_{c}$ on both sides arises solely from the finite infrared extent of the slow–mode reservoir, without invoking competing orders or fine tuning. (b) Uemura scaling collapse between superconducting transition temperature and superfluid stiffness, $T_{c}\propto\rho_{s}$. Both pairing strength and phase coherence are controlled by the same infrared spectral weight $\rho_{0}$ of slow collective modes, yielding universal linear scaling across material families.
@@
The linearized transition scale derived in Eq. (60) shows that superconductivity is governed by the infrared spectral weight $\rho_{0}$ of the slow–mode reservoir. We interpret this regime not as a finely tuned quantum critical point, but as a self–organized dynamical fixed point of the relaxation spectrum.
In this phase, feedback between electronic correlations and collective relaxation continuously regenerates an extensive near–marginal sector, leading to a scale–free infrared TDOS,
$\rho(\lambda\to 0)=\rho_{*}>0,$ (62)
without the need for fine tuning of microscopic parameters. The flat infrared spectrum therefore characterizes a dynamical phase rather than a measure–zero critical point.
External control parameters such as doping or pressure primarily modulate the effective infrared cutoff scale of this reservoir. Let $p$ denote a tuning parameter and $p_{\star}$ the point at which the infrared cutoff is minimal. The collective mass term behaves as
$r(p)\sim|p-p_{\star}|,$ (63)
so that the associated cutoff obeys
@@
Away from $p_{\star}$, the slow–mode reservoir remains present but is truncated in the deep infrared, yielding
@@
with $f$ a decreasing crossover function.
Substituting into Eq. (60) gives
$T_{c}(p)\sim E_{\rm pair}\,\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (66)
which naturally produces a superconducting dome. Importantly, the suppression of $T_{c}$ away from $p_{\star}$ does not arise from the disappearance of the dynamical fixed point, but from the progressive truncation of its infrared extent.
In this interpretation, optimal doping does not correspond to a conventional symmetry–breaking quantum critical point. Rather, it marks the regime in which the self–organized infrared spectral phase is most fully exposed.
@@
While Sec. IV established that the pairing scale is algebraically enhanced by the infrared TDOS, the observable transition temperature is ultimately limited by global phase coherence. In systems with reduced superfluid density, the ordering temperature may be controlled by phase stiffness rather than by gap formation.
@@
Within the slow–mode reservoir framework, the same infrared spectral weight that enhances pairing also governs coherent charge transport. The superfluid stiffness measures the energy cost of imposing a spatial phase twist, and is therefore proportional to the density of coherent paired carriers that remain dynamically correlated over long time scales.
In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$. Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$, the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight. Integrating out amplitude fluctuations in the presence of the slow-mode reservoir generically yields
$\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (69)
up to nonuniversal numerical factors.
Both pairing amplitude and phase rigidity are thus controlled by a single infrared parameter that quantifies the extent of the slow–mode reservoir.
Combining Eqs. (68) and (69) gives
$T_{c}\sim\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (70)
establishing a direct proportionality between the superconducting transition temperature and superfluid density.
This relation reproduces the empirical Uemura scaling observed across cuprates and other unconventional superconductors, where $T_{c}$ tracks the zero-temperature superfluid density. In contrast to weak-coupling BCS superconductors—where phase stiffness is large and essentially decoupled from gap formation—correlated systems operate in a regime in which infrared collective dynamics simultaneously controls pairing and coherence.
Within the present framework, this proportionality is not imposed phenomenologically. It follows directly from the TDOS-controlled infrared fixed point, where a single spectral parameter $\rho_{0}$ governs both the Cooper instability and phase stiffness. The observed scaling of $T_{c}$ with superfluid density thus emerges as a natural consequence of memory-dominated critical dynamics.
@@
Such long-time tails naturally produce 1/f-type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors. Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure. Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization.
V.4 D. Unified infrared scaling picture
We emphasize that all experimental signatures discussed above—the dome, Uemura scaling, and long–time dynamical correlations—originate from the same infrared parameter $\rho_{0}$, the low–energy spectral weight of the slow–mode reservoir.
The density of relaxation rates controls: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response. Quantum criticality thus acts not merely as a source of fluctuations but as a dynamical amplifier governed by infrared spectral reorganization that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) are expected to display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling. The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix E.
Figure 4: Dynamical classification determined by the TDOS exponent $\alpha$. The scaling $\rho(\lambda)\sim\lambda^{\alpha}$ implies $\Pi(\omega)\sim|\omega|^{\alpha-1}$ and separates the system into memory–dominated ($\alpha<1$), marginal Ohmic ($\alpha=1$), and analytic ($\alpha>1$) regimes, with corresponding pairing consequences. Here $\omega$ denotes Matsubara frequency.
@@
The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum rather than solely by the softening of a single order parameter. The time–scale density of states provides a quantitative measure of this organization. When the TDOS remains finite as $\lambda\to 0$, the system enters a memory–dominated regime characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions.
@@
This viewpoint clarifies the limits of conventional coarse-grained field theories. Away from criticality, the relaxation spectrum is gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective fields. Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending toward zero. In this regime internal modes cannot be integrated out: the full decay spectrum remains dynamically active across time scales. The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a universal consequence of dynamical criticality.
@@
This observation highlights a structural difference from Hertz–Millis–type approaches. Conventional quantum critical theories assume overdamped local dynamics of a small set of bosonic modes, with dissipation controlled by Landau damping. In contrast, the present framework allows for spectral condensation of near-marginal relaxation channels. Criticality is reinterpreted not as single-mode softening but as a reorganization of the entire decay spectrum. The TDOS therefore replaces phenomenological damping coefficients with a microscopic spectral quantity governing infrared universality.
@@
This structure is summarized schematically in Fig. 4. The figure illustrates how the infrared universality class of the collective dynamics is determined by the exponent of the relaxation-rate density of states. For a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$, the polarization kernel assumes the universal infrared form $\Pi(\omega)\sim|\omega|^{\alpha-1}$. Different dynamical regimes therefore emerge as special cases of the same spectral organization. For $\alpha>1$ the density of slow modes vanishes in the infrared, so that the polarization kernel remains finite and the collective dynamics approaches the weakly damped limit, with the infrared response dominated by the nondissipative part $\chi^{-1}\simeq r+cq^{2}$. The marginal case $\alpha=1$ corresponds to Ohmic dissipation, reproducing the Hertz–Millis form $\chi^{-1}\simeq r+cq^{2}+\gamma|\omega|$. Most importantly, when the TDOS approaches a finite constant ($\alpha=0$), a dense continuum of near-marginal relaxation modes emerges. In this regime the polarization acquires the singular form $\Pi(\omega)\sim 1/|\omega|$, producing long-time memory kernels $K(t)\sim 1/t$ and the non-Markovian infrared dynamics discussed throughout this work.

### 30 Mar 2026 v5

@@ Abstract
the retarded kernel exhibits logarithmic scaling,
$\mathrm{Re}\,\Pi^{R}(\omega)\sim\rho_{0}\ln\frac{\Lambda}{|\omega|}.$ (50)
Substituting $\omega\sim T_{c}$ into Eq. (33) yields
$1-g_{\rm pair}\,\rho_{0}\ln\frac{\Lambda}{T_{c}}=0,$ (51)
from which the transition temperature follows as
$T_{c}\sim\Lambda\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}}\right).$ (52)
Thus, a flat TDOS leads to a marginal enhancement of pairing, reproducing a BCS-like exponential transition scale despite the presence of strong infrared fluctuations.
Infrared-divergent TDOS ($\alpha<0$): IR-singular regime. When the relaxation spectrum develops an infrared divergence,
$\rho(\lambda)\sim\lambda^{\alpha},\qquad\alpha<0,$ (53)
the real part of the retarded kernel acquires a power-law infrared singularity,
$\mathrm{Re}\,\Pi^{R}(\omega)\sim C_{\alpha}|\omega|^{\alpha}.$ (54)
At finite temperature, thermal fluctuations cut off the infrared divergence at the scale $\lambda_{\rm IR}\sim T$, so that the static retarded kernel should be evaluated with the effective lower limit set by $T$:
$\Pi^{R}(0;T)\sim\int_{T}^{\Lambda}d\lambda\;\frac{\rho(\lambda)}{\lambda}.$ (55)
For an infrared-singular TDOS, this gives
$\Pi^{R}(0;T)\sim\int_{T}^{\Lambda}d\lambda\;\lambda^{\alpha-1}=\frac{T^{\alpha}-\Lambda^{\alpha}}{|\alpha|}\sim\frac{1}{|\alpha|}T^{-|\alpha|}.$ (56)
The Thouless criterion therefore becomes
$1-g_{\rm pair}C_{\alpha}T_{c}^{-|\alpha|}=0,$ (57)
yielding
$T_{c}\sim(g_{\rm pair}C_{\alpha})^{1/|\alpha|}\sim g_{\rm pair}^{1/|\alpha|}.$ (58)
The enhancement becomes stronger for larger $|\alpha|$, since the retarded kernel diverges more strongly in the infrared as $T^{-|\alpha|}$.
In this regime, the accumulation of slow modes toward vanishing relaxation rates reorganizes the Cooper channel into a relevant infrared interaction. As a result, the pairing instability is governed by a power-law infrared enhancement of the retarded kernel, and the exponential suppression characteristic of conventional pairing is replaced by an algebraic dependence on the interaction strength.
Physical interpretation. The transition temperature derived above should be understood as the pairing instability scale determined by the retarded Cooper kernel. It is therefore controlled by the infrared spectral structure of the relaxation modes. A finite TDOS produces strong fluctuations but only logarithmic enhancement of the retarded kernel, leading to an exponentially moderate pairing scale. By contrast, an infrared-divergent TDOS modifies the retarded response itself, generating algebraic growth of the pairing interaction and a parametrically enhanced instability scale.
In real materials, however, the observable superconducting transition requires not only the formation of pairs but also global phase coherence. The actual transition temperature is therefore bounded by the smaller of the pairing scale and the superfluid stiffness,
@@
with $\rho_{s}\sim n_{s}/m^{*}$. The scaling of phase stiffness within the present framework is discussed in Sec. V.
The central result is that the pairing instability itself is governed by a single infrared characteristic of the system, namely the low-energy structure of the TDOS. Superconductivity thus emerges as a consequence of dynamical spectral organization, providing an alternative to descriptions based solely on the exchange of specific bosonic mediators.
Figure 3: Experimental consequences of memory–dominated critical dynamics. (a) Schematic superconducting dome emerging from the infrared spectral structure of the slow–mode reservoir. The transition temperature follows $T_{c}(p)\sim\Lambda(p)\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}(p)}\right)$ and is maximal near $p_{\star}$, where both the infrared spectral weight and the extent of the slow–mode reservoir are largest. Suppression of $T_{c}$ away from this regime arises from the combined effect of reduced infrared spectral weight and finite infrared cutoff, without invoking competing orders. (b) Uemura scaling showing the proportionality between superconducting transition temperature and superfluid stiffness, $T_{c}\propto\rho_{s}$, in regimes where phase coherence limits the transition. Both pairing and phase stiffness are influenced by the same infrared spectral weight $\rho_{0}$, providing a common dynamical origin for the observed scaling across material families.
@@
[…]
We interpret this regime not as a finely tuned quantum critical point, but as a self–organized dynamical state characterized by an extended distribution of slow relaxation modes. In this state, feedback between electronic correlations and collective dynamics maintains a broad near-marginal sector of the TDOS,
$\rho(\lambda\to 0)=\rho_{*}(p),$ (60)
whose magnitude and infrared extent depend on external control parameters such as doping or pressure.
Let $p$ denote a tuning parameter and $p_{\star}$ the point at which the infrared extent of the slow–mode reservoir is maximal. Deviations from $p_{\star}$ introduce a finite mass scale $r(p)$ that truncates the infrared spectrum,
@@
At the same time, the low-energy spectral weight evolves smoothly,
@@
where $f$ is a decreasing crossover function.
Substituting these dependencies into Eq. (52) yields
$T_{c}(p)\sim\Lambda(p)\,\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}(p)}\right),$ (63)
which naturally produces a superconducting dome.
The physical origin of the dome can be understood as a competition between two infrared effects. Near $p_{\star}$, the slow–mode reservoir is maximally extended, leading to a large $\rho_{0}$ and a large cutoff $\Lambda$, and hence enhanced pairing. Away from this regime, superconductivity is suppressed either by the reduction of the infrared spectral weight (overdoped side) or by the progressive truncation of the slow–mode reservoir (underdoped side).
In this interpretation, optimal doping does not correspond to a fine-tuned symmetry-breaking quantum critical point. Rather, it marks the regime in which the infrared spectral organization of relaxation modes is most fully developed. The superconducting dome therefore emerges as a robust consequence of infrared spectral reorganization, without requiring a specific bosonic mediator or fine tuning of microscopic interactions.
@@
While Sec. IV establishes that the pairing instability is governed by the infrared structure of the retarded Cooper kernel, the observable transition temperature is ultimately constrained by global phase coherence. In systems with reduced carrier density or strong correlations, the superconducting transition may therefore be limited by phase stiffness rather than by pair formation.
@@
Within the present framework, both pairing and phase coherence are influenced by the same infrared spectral structure of the relaxation modes. The superfluid stiffness measures the energy cost of imposing a spatial phase twist, and is therefore controlled by the density of carriers that remain phase coherent over long time scales. In a memory–dominated regime, such long-time coherence is governed by the low-energy spectral weight $\rho_{0}$ of the slow–mode reservoir.
As a result, the superfluid stiffness is expected to scale with the infrared spectral weight,
$\rho_{s}\sim\rho_{0},$ (66)
up to material-dependent prefactors that depend on microscopic details such as band structure and effective mass.
Combining Eqs. (65) and (66), one expects that the superconducting transition temperature tracks the superfluid density,
$T_{c}\propto\rho_{s},$ (67)
consistent with the empirical Uemura scaling observed in cuprates and other unconventional superconductors.
In contrast to weak-coupling BCS superconductors, where phase stiffness is large and essentially decoupled from gap formation, strongly correlated systems operate in a regime where both pairing and coherence are governed by infrared collective dynamics. Importantly, however, the transition temperature is determined by the smaller of the pairing scale and the phase stiffness, $k_{B}T_{c}\sim\min\left(\Delta,\,\rho_{s}\right)$, so that Uemura scaling emerges most clearly in regimes where phase stiffness provides the dominant limitation.
Within this picture, the observed proportionality between $T_{c}$ and superfluid density does not arise from a direct identity between pairing strength and phase rigidity, but from their common dependence on the infrared spectral weight of the slow–mode reservoir. The TDOS therefore acts as a unifying control parameter that links pairing susceptibility and phase coherence without requiring a specific microscopic pairing mechanism.
@@
Such long-time tails naturally produce 1/f-type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors. Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure. Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization. Related derivations are presented in Appendix E.
Unified infrared scaling picture. We emphasize that all experimental signatures discussed above—the dome, Uemura scaling, and long–time dynamical correlations—originate from the same infrared parameter $\rho_{0}$, the low–energy spectral weight of the slow–mode reservoir.
The density of relaxation rates plays a central organizing role: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response. Quantum criticality thus acts not merely as a source of fluctuations but as a dynamical amplifier governed by infrared spectral reorganization that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) are expected to display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling. The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix F.
Figure 4: Dynamical classification determined by the TDOS exponent $\alpha$. The scaling $\rho(\lambda)\sim\lambda^{\alpha}$ implies $\Pi^{R}(\omega)\sim|\omega|^{\alpha}$ (with logarithmic behavior at $\alpha=0$), and separates the system into infrared-singular ($\alpha<0$), marginal ($\alpha=0$), Ohmic ($\alpha=1$), and analytic ($\alpha>1$) regimes, with corresponding pairing consequences. Here $\omega$ denotes frequency (Matsubara or real, depending on context).
@@
The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum, rather than being described solely by the softening of a single order parameter. The time–scale density of states provides a quantitative measure of this organization. When the TDOS remains finite as $\lambda\to 0$, the system enters a memory–dominated regime characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions.
@@
This viewpoint clarifies the limits of conventional coarse-grained field theories. Away from criticality, the relaxation spectrum is gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective fields. Near correlated critical regimes, however, relaxation rates form a dense continuum extending toward zero. In this regime, internal modes cannot be trivially integrated out: the decay spectrum remains dynamically relevant across a wide range of time scales. The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a generic consequence of dynamical criticality.
@@
This observation highlights a structural difference from Hertz–Millis–type approaches. Conventional quantum critical theories assume overdamped local dynamics of a small set of bosonic modes, with dissipation controlled by Landau damping. In contrast, the present framework allows for spectral condensation of near-marginal relaxation channels. Criticality is reinterpreted not as single-mode softening but as a reorganization of the decay spectrum. The TDOS therefore provides a microscopic spectral quantity governing infrared universality, complementing phenomenological damping descriptions.
@@
This structure is summarized schematically in Fig. 4. The figure illustrates how the infrared universality class of the collective dynamics is determined by the exponent of the relaxation-rate density of states. For a power-law TDOS, $\rho(\lambda)\sim\lambda^{\alpha}$, the retarded polarization kernel exhibits the infrared scaling $\mathrm{Re}\,\Pi^{R}(\omega)\sim|\omega|^{\alpha}$ for $\alpha<1$, with logarithmic behavior in the marginal case $\alpha=0$. The inverse susceptibility is correspondingly given by $\chi^{-1}(q,\omega)\simeq r+cq^{2}+\Pi^{R}(\omega)$, so that different dynamical regimes emerge as special cases of a unified spectral organization. For $\alpha>1$, the density of slow modes vanishes in the infrared, and the dynamical correction becomes subleading compared to the static terms at low frequencies. In this regime the collective dynamics approaches a weakly damped, regular (analytic) limit, with the infrared response dominated by the nondissipative contribution. The case $\alpha=1$ corresponds to the Ohmic boundary. In this case the kernel acquires a linear frequency dependence up to logarithmic corrections, $\Pi^{R}(\omega)\sim|\omega|$, reproducing the Hertz–Millis form $\chi^{-1}\simeq r+cq^{2}+\gamma|\omega|$. Most importantly, when the TDOS approaches a finite constant ($\alpha=0$), a dense continuum of near-marginal relaxation modes emerges. In this regime the polarization exhibits logarithmic infrared behavior, $\mathrm{Re}\,\Pi^{R}(\omega)\sim\rho_{0}\ln(\Lambda/|\omega|)$, while the absorptive part remains finite at low frequencies. This leads to long-time memory kernels $K(t)\sim 1/t$ and the non-Markovian infrared dynamics discussed throughout this work.